### v1

# GRPO-LEAD: A Difficulty-Aware Reinforcement Learning Approach for Concise Mathematical Reasoning in Language Models

Abstract
Recent advances in R1-like reasoning models leveraging Group Relative Policy Optimization (GRPO) have significantly improved the performance of language models on mathematical reasoning tasks. However, current GRPO implementations encounter critical challenges, including reward sparsity due to binary accuracy metrics, limited incentives for conciseness, and insufficient focus on complex reasoning tasks. To address these issues, we propose GRPO-LEAD, a suite of novel enhancements tailored for mathematical reasoning. Specifically, GRPO-LEAD introduces (1) a length-dependent accuracy reward to encourage concise and precise solutions, (2) an explicit penalty mechanism for incorrect answers to sharpen decision boundaries, and (3) a difficulty-aware advantage reweighting strategy that amplifies learning signals for challenging problems. Furthermore, we systematically examine the impact of model scale and supervised fine-tuning (SFT) strategies, demonstrating that larger-scale base models and carefully curated datasets significantly enhance reinforcement learning effectiveness. Extensive empirical evaluations and ablation studies confirm that GRPO-LEAD substantially mitigates previous shortcomings, resulting in language models that produce more concise, accurate, and robust reasoning across diverse mathematical tasks. Our source code, generated dataset, and model are available at https://github.com/aeroplanepaper/GRPO-LEAD.
1 Introduction
Figure 1: Validation Pass@1 over training steps for three configurations: GRPO, GRPO with length reward, and GRPO with length reward plus advantage reweighting. The validation consists of 27 challenging problems from AIMO2 Frieder et al. (2024), CMU-MATH-AIMO’s validation Sun (2024), and AIME24, where Deepseek-14B struggles to solve.
Recently, R1-like reasoning models have attracted significant attention due to their impressive performance in solving challenging mathematical reasoning tasks through extensive chains of thoughtLuo et al. (2025); Wen et al. (2025). According to the technical report introducing R1Guo et al. (2025), reinforcement learning (RL) fine-tuning plays a pivotal role in enabling this reasoning capability. In particular, Group Relative Policy Optimization (GRPO)Shao et al. (2024), a novel RL approach for language models, has emerged as a promising alternative to traditional methods such as PPOSchulman et al. (2017) and DPORafailov et al. (2023), primarily due to its efficiency and intrinsic compatibility with language model training.
However, existing GRPO implementations still encounter substantial limitations. One key issue is the inherent reward sparsity arising from binary and rule-based accuracy metrics, which significantly hampers effective model training. Specifically, when all generated responses to a given problem are either uniformly correct or incorrect, the resulting uniform reward signal provides minimal differentiation, leading to weak learning gradients and consequently slower model convergence. To elaborate, if all outputs in a question group are correct, each receives identical positive feedback, diluting the informative gradient needed for meaningful policy improvements. Conversely, uniformly incorrect responses yield no useful information to guide policy refinement.
Furthermore, computational efficiency also emerges as a critical practical concern, as reinforcement learning fine-tuning typically demands substantial resources, limiting accessibility, experimentation speed, and scalability, especially in low-resource environments. The current GRPO formulation is insufficient for encouraging concise and precise reasoning. Consequently, reducing computational requirements during both training and inference is essential for enabling broader applicability and effective real-world deployment.
Motivated by these limitations, this work introduces GRPO-LEAD, a suite of targeted modifications explicitly designed to enhance GRPO’s effectiveness for mathematical reasoning tasks. Our key contributions include:
  1. 1.
    Length-dependent Accuracy Reward: Introduces a dynamic reward shaping mechanism that promotes brevity among correct responses using standardized length-based penalties, reducing verbosity without sacrificing accuracy.
  2. 2.
    Explicit Penalty for Incorrect Solutions: Implements a negative reward for incorrect outputs to enforce a sharper decision boundary, mitigating overconfidence and boosting precision of the model.
  3. 3.
    Difficulty-aware Advantage Reweighting: Applies a logistic weighting function to advantage estimates based on empirical correctness rates, ensuring stronger updates for harder problems and balanced generalization.
  4. 4.
    Impact of Model Scale and Data Quality on Reinforcement Learning Effectiveness: Demonstrates that larger base models and high-quality, curriculum-structured fine-tuning data significantly improve RL convergence and output quality; also introduces targeted reward interventions to mitigate repetitive formatting artifacts.
Empirical evaluations and comprehensive ablation studies confirm that our method effectively addresses previous GRPO shortcomings, leading to more concise, accurate, and efficiently trainable models capable of robust performance across mathematical reasoning tasks.
2 Related Work
2.1 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO), a recently proposed algorithm designed specifically for fine-tuning language models with group-level normalization of rewardsGuo et al. (2025). GRPO modifies the standard policy gradient objective by introducing a relative advantage within sets of responses corresponding to the same question, stabilizing updates, and promoting consistent learning signals. Formally, GRPO defines the objective as:
$\mathcal{L}_{\text{GRPO}}(\theta)=-\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Bigg{[}$ $\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}\hat{A}_{i,t}$
$\quad-\beta D_{\text{KL}}[\pi_{\theta}\|\pi_{\theta_{\text{old}}}]\Bigg{]},$ (1)
where $G$ is the number of groups, $\hat{A}_{i,t}$ is the normalized advantage within the group, and $\beta$ controls the KL divergence penalty term enforcing policy stability.
2.2 Length Reward
A prevalent issue in reinforcement learning-based fine-tuning of language models is the uncontrolled increase in response length driven by reward signals, commonly known as reward hacking. This phenomenon leads to unnecessarily verbose responses, which, although technically correct, often lack conciseness and hinder interpretability. Furthermore, such verbosity fails to reflect efficient reasoning, limiting model utility in practical scenarios. Existing efforts to mitigate this problem typically involve incentivizing shorter answers to encourage more succinct reasoning processes. For example, Kimi proposed an individual min-max normalized length reward based on the lengths of generated responses Team et al. (2025). Yeo et al. introduced a cosine length reward function with fixed maximum and minimum thresholds to manage response lengths Yeo et al. (2025). Aggarwal et al. utilized a target "golden length" to directly reward or penalize responses based on their deviation from an ideal length Aggarwal and Welleck (2025).
However, these existing methods depend heavily on static or predefined length heuristics, limiting their effectiveness across diverse questions of varying complexity. In contrast, our proposed length-dependent accuracy reward addresses these limitations by dynamically calibrating rewards according to each group’s relative response length and rollout accuracy, promoting concise yet difficulty-aware reasoning processes.
3 Method
To systematically address the limitations identified in existing implementations of Group Relative Policy Optimization (GRPO), we propose a suite of novel modifications collectively termed GRPO-LEAD (GRPO with Length-dependent rewards, Explicit penalties, and Advantage reweighting for Difficulty). Our proposed method enhances the original GRPO framework by introducing three core innovations: 1) a length-dependent accuracy reward to foster concise solutions, 2) an explicit penalty mechanism to mitigate low precision rate caused by length reward, and 3) an advantage reweighting strategy tailored to compensate for limited rollout scenarios. Additionally, we rigorously examine how base model scale and supervised fine-tuning (SFT) impact the effectiveness of reinforcement learning (RL) fine-tuning.
3.1 Length-Dependent Accuracy Reward
The core idea is to reward correct completions not uniformly but in proportion to their relative conciseness. Given a question $q$ and a set of model-generated responses $\{o_{i}\}$, we first isolate the subset of correct responses and compute the mean $\mu$ and standard deviation $\sigma$ of their token lengths. For a correct response $o$ with length $|o|$, we define its standardized length deviation as:
$z=\frac{|o|-\mu}{\sigma+\epsilon},$ (2)
where $\epsilon>0$ is a small constant added for numerical stability. The final reward is modulated using an exponential decay function:
$R_{\text{accuracy}}(o|q)=\begin{cases}\exp(-\alpha z),&\text{if }o\text{ is correct},\\ 0,&\text{if }o\text{ is incorrect}.\end{cases}$ (3)
where $\alpha>0$ is a tunable hyperparameter controlling the strength of length penalization.
This formulation ensures that overly long correct responses are systematically penalized, while relatively concise ones are amplified. Unlike static or absolute length constraints, our approach leverages standardized deviation, allowing for dynamic adaptation to the distributional properties of each question.
3.2 Explicit Penalty for Incorrect Answers
Existing methods often focus on maximizing pass@1 within a restricted context length. However, we observe that prioritizing pass@1 can reduce overall accuracy. Our experiments indicate that this decline is not directly caused by incorporating a length reward but rather by relying on a binary accuracy reward.
When length-based regularization is applied, it may shorten responses and limit the model’s ability to “re-think” or revise incorrect answers through longer reasoning. Still, disabling the length reward does not fully address the problem. Although pass@1 improves for both training and evaluation, the solve-all metric—measured by how many questions are answered correctly in all sampled responses—tends to drop. Adding the length reward simply accelerates this trend.
We attribute this outcome to the binary reward signal. Under a limited generation length, a model is incentivized to complete its response within the available budget. While such an approach may be less rigorous and more prone to errors, it can still yield a non-zero reward, since providing a guess is better than providing none at all. This creates a shortcut, where ambiguous or partially complete answers inflate pass@1 yet lower overall precision.
To mitigate this, we propose a revised reward structure that explicitly penalizes incorrect responses, thereby reinforcing a sharper decision boundary between correct and incorrect outputs. The revised reward function is defined as:
$R_{\text{accuracy}}(o\mid q)=\begin{cases}\exp(-\alpha z),&\text{if }o\text{ is correct},\\ -1,&\text{if }o\text{ is incorrect},\end{cases}$ (4)
where $z$ denotes the standardized length deviation, and $\alpha>0$ controls the strength of length penalization for correct responses, as previously defined.
Under this formulation, the expected reward for a response with correctness probability $P(\text{correct})$ is:
$\mathbb{E}[R_{\text{accuracy}}(o\mid q)]$ $=P(\text{correct})\cdot\exp(-\alpha z)$
$\quad-(1-P(\text{correct}))$ (4)
To gain intuition about the behavior of this reward function, we consider a simplified case where the length penalty is neutralized, i.e., $\exp(-\alpha z)\approx 1$(In practice the mean reward of correct answer is approcximately 1). Under this assumption:
$\mathbb{E}[R]\approx 2P(\text{correct})-1$ (5)
This approximation highlights a key property of the reward: the expected value becomes positive only when $P(\text{correct})>0.5$.
This formulation introduces a principled deterrent against random guessing and encourages the model to internalize a more robust decision threshold. Empirically, this approach improves both pass@1 and overall precision.
3.3 Advantage Reweighting for Difficulty-Aware Training
While the length-dependent accuracy reward and explicit penalty formulation described earlier significantly mitigate verbosity and enhance precision, a subtle yet critical challenge remains. In the original GRPO algorithm, the rewards are uniformly applied across questions regardless of intrinsic difficulty. This uniformity implicitly biases the model to excessively optimize performance on simpler tasks—where obtaining correct and concise responses is inherently easier—while inadvertently neglecting harder questions that require deeper reasoning and longer deliberation. Consequently, the model may degrade on complex problems due to an insufficient incentive structure.
To address this imbalance, we introduce a difficulty-aware advantage reweighting strategy designed to dynamically adjust the magnitude of policy updates based on problem difficulty. Intuitively, we aim to amplify learning signals for harder problems, thereby encouraging sustained exploration and deeper reasoning strategies rather than superficial optimization on trivial tasks.
Formally, given a prompt set grouped by question $q$ with associated responses ${o_{i}}$, we first define the group’s empirical correctness ratio:
$\rho_{q}=\frac{\text{number of correct responses for }q}{\text{total number ofresponses for }q}.$ (5)
This correctness ratio serves as a proxy for problem difficulty, with lower values indicating greater difficulty.
Next, we introduce a logistic reweighting factor dependent on this ratio to modulate the advantage estimates during the RL training step. The logistic function is defined as:
$w(\rho_{q})=A+\frac{B-A}{1+\exp\left[k(\rho_{q}-\rho_{0})\right]},$ (6)
where hyperparameters $A,B,\rho_{0},k$ allow precise control over the sensitivity of weighting to problem difficulty.
To apply this weighting to the advantage calculation, consider the normalized advantage estimate for response $o_{i}$:
$\tilde{A}i=\frac{R(o_{i}|q)-\mu_{q}}{\sigma_{q}+\epsilon},$ (7)
where $\mu_{q}$ and $\sigma_{q}$ represent the mean and standard deviation of rewards within the prompt group $q$. We then define the difficulty-aware advantage as:
$A_{i}^{\prime}=\tilde{A}_{i}\cdot\begin{cases}w(\rho_{q}),&\text{if }\tilde{A}_{i}>0\\ w(1-\rho_{q}),&\text{if }\tilde{A}_{i}\leq 0\end{cases}$ (8)
This formulation ensures that for difficult problems (low $\rho_{q}$), correct responses (which are rare and thus highly valuable) receive substantially larger updates due to the increased weighting $w(\rho_{q})$. Conversely, incorrect responses on easier problems (high $\rho_{q}$) are penalized more strongly, sharpening the decision boundary for problems where high performance should be expected.
3.4 Impact of Model Scale and Data Quality on Reinforcement Learning Effectiveness
While the preceding reward strategies effectively address verbosity and improve precision across varying difficulties, the base model’s scale and the quality of the underlying dataset significantly influence the effectiveness of RL process. Specifically, we analyze how initial model capability and carefully curated training data jointly affect RL fine-tuning outcomes.
Firstly, our experiments reveal a clear dependence between model scale and RL improvements. While RL fine-tuning notably enhances the Deepseek Distilled 7B model’s performance on relatively straightforward questions, substantial gains on complex reasoning tasks remain elusive. The 7B model often prematurely converges to incorrect reasoning paths and frequently overlooks critical edge cases. Conversely, we observe that the Deepseek Distilled 14B model inherently outperforms even the RL-enhanced 7B variant, particularly on questions requiring exhaustive reasoning or enumeration of multiple scenarios. These findings suggest intrinsic limitations in smaller models’ capacities for advanced reasoning, underscoring the importance of larger-scale base models for RL fine-tuning to effectively tackle complex problems.
To further investigate methods for enhancing model capability, we generated a targeted dataset comprising 13k math reasoning problems sourced from the dataset provided by DeepScalerLuo et al. (2025), including historical AMC, AIME, and OmniMath exams, accompanied by solutions generated using QwQ32BTeam (2025), a state-of-the-art small scale reasoning model. After supervised fine-tuning (SFT) with this specialized dataset, we applied our proposed RL strategies. Despite initial signs of overfitting in the SFT stage, we observed that subsequent RL fine-tuning rapidly alleviated these issues, demonstrating faster convergence and significant improvements in both pass@1 accuracy and overall precision relative to RL fine-tuning applied directly to the original model.
Our results additionally highlight the critical role of data quality and curriculum strategies in sustained RL improvement. By first applying RL on a subset of approximately 7k challenging problems (difficulty rating $\geq 4$ from the DeepScaler dataset), we obtained a robust initial policy checkpoint. Subsequently, we refined this policy using a curriculum consisting of the most challenging problems identified from the first-stage correctness distribution (correctness rate below 50%), supplemented by high-difficulty examples from the second stagen dataset of Light-R1 Wen et al. (2025). Empirical evaluations demonstrate that this two-stage curriculum significantly enhances the model’s ability to continuously improve on complex tasks.
Finally, we addressed a persistent formatting issue—repetitive n-gram patterns likely arising from the absence of clear end-of-sequence (EOS) signals during the initial SFT stage. By temporarily removing length-dependent rewards and introducing an explicit negative reward ($-1.5$) for repeated n-grams, we achieved further improvements in precision and pass@1 metrics. This intervention demonstrates the effectiveness of targeted reward modifications in mitigating specific output-formatting anomalies.
In summary, our experiments affirm the substantial influence of initial model capacity, dataset quality, and targeted reward engineering on RL-based fine-tuning outcomes. These findings collectively inform a systematic approach for enhancing language models’ capability to produce concise, accurate, and well-structured responses across tasks of varying complexity.
4 Experimental Setup
We evaluate GRPO-LEAD, integrating length-dependent accuracy rewards, explicit penalties for incorrect solutions, and difficulty-aware advantage reweighting, on DeepSeek-R1 Distilled variants Guo et al. (2025); Yang et al. (2024). Our experiments cover two model scales, 7B and 14B parameters. All GRPO training is conducted using the VERL framework.Sheng et al. (2024).
4.1 Datasets and Filtering
Our primary training data is sourced from the DeepScaler dataset Luo et al. (2025). We begin by filtering out problems with difficulty ratings below 2.5, resulting in approximately 9,000 questions for fine-tuning.
For stages 2 and 3 of our 14B model experiments, we further refine the dataset by selecting problems where the model’s stage-1 rollout accuracy is no greater than 75%, yielding around 2,283 questions. Additionally, we incorporate challenging problems with numeric answers from the stage-2 dataset of Light-R1 Wen et al. (2025).
In total, the dataset for stages 2 and 3 comprises 3,524 questions. This adaptive filtering strategy ensures a focused emphasis on harder problems, aiming to improve the model’s performance on more complex tasks.
4.2 Hyperparameters
We train with a learning rate of $1\times 10^{-6}$, batch size 32, and group size 8–generating 8 rollouts per question for GRPO reward computation. The KL penalty term is removed, as it was found to suppress exploration in our experiments, which is also suggested in similar worksLiu et al. (2025); Hu et al. (2025).
For the length-dependent accuracy reward, we set $\alpha=0.05$, providing a moderate decay that encourages conciseness without penalizing slight verbosity.
For difficulty-aware advantage reweighting, we use $A=0.4$, $B=1.5$, $\rho_{0}=0.75$, and $k=10$. This configuration ensures reweighting is minimal on easy problems but sharply increases near the 75% correctness threshold. The steep slope ($k=10$) enables strong emphasis on high-difficulty examples, guiding the model to allocate learning more effectively.
4.3 Model Variants and Fine-Tuning Stages
7B Model Experiments
Starting from the DeepSeek-R1 Distilled 7B Qwen-Math checkpoint, we first apply standard GRPO on the 9k questions, producing a baseline. Then, we train 3 more models from the DeepSeek-R1 Distilled 7B Qwen-Math checkpoint, adding one more of the following components subsequently: (i) Length Reward only, (ii) Length Reward + Advantage Reweighting, (iii) Length Reward + Advantage Reweighting + Explicit Penalty. We train for approximately 200 steps and select the top-performing checkpoints based on validation results. At test time, we limit the generation length to 8k for all 7B models, matching the training length limit.
14B Model Experiments
We extend the above procedure to the DeepSeek-R1 Distilled 14B Qwen checkpoint across multiple stages. In Stage 1, we train for 100 steps using all GRPO-LEAD components on the filtered 9k-question dataset. To evaluate the benefit of supervised fine-tuning (SFT), we first fine-tune the model on a curated set of 13k math problems, then apply GRPO. This SFT stage significantly improves the model’s reasoning quality, even though it tends to increase the output length and caused some format error.
The supervised fine-tuning data consists of all problems in the DeepScaler dataset with difficulty greater than 1. To construct high-quality reasoning traces for SFT, we use the QWQ-32B modelTeam (2025) to generate step-by-step solutions.
After observing that some questions remain low correctness, we further fine-tune for Stage 2 with an additional 100 steps, focusing on these underperformed problems. Lastly, in Stage 3, we address repetitive output patterns by removing the length penalty and introducing a negative reward ($-1.5$) for repeated $n$-grams. We continue training for 140 more steps, yielding the final model checkpoint. At test time, we limit the generation length to 14k for all 14B models, in accordance with our training settings and also to better evaluate the models’ performance in a low-budget scenario.
4.4 Baselines and Evaluation Protocol
We compare our models with both DeepSeek-R1 Distilled-14B-QwenGuo et al. (2025) (the distilled Qwen model without GRPO-LEAD) and Light-R1-14B-DS Wen et al. (2025), which has the same base model as ours and was first finetuned with 3k hard math problems with SFT, and then fine-tuned with a cosine-based length reward Yeo et al. (2025) on their selected math problems for three epochs using GRPO.
We primarily report three metrics: (1) Cons@32, accuracy through majority voting for 32 samplings; (2) Pass@1, the probability that the top-1 sample is correct under a chosen decoding strategy; (3) Average Length (Lenavg), measuring verbosity. Unless otherwise specified, we decode with temperature 0.6 and sample 32 solutions per question, then compute Cons@32 and Pass@1 over these samples.
Table 1: Ablation results on AIME24 and AIME25. We report Cons@32 (the fraction of problems for which at least one correct solution is found among 32 samples), Pass@1, and the average token length (Lenavg). The best value in each column is in boldface, the second best is underlined.
Ablation Setting AIME24 AIME25
Cons@32 Pass@1 Lenavg Cons@32 Pass@1 Lenavg
Deepseek-7B 0.767 0.431 6990 0.467 0.292 7113
GRPO + len. reward 0.767 0.438 5275 0.533 0.308 5210
     + adv. reweighting 0.767 0.458 5323 0.567 0.325 5437
     + explicit penalty 0.800 0.470 6104 0.567 0.345 6308
Table 2: Comparison of model performance on AIME24 and AIME25, showing Cons@32, Pass@1, and average token length (Lenavg). The best value in each column is in boldface, the second best is underlined.
Model Name AIME24 AIME25
Cons@32 Pass@1 Lenavg Cons@32 Pass@1 Lenavg
DeepSeek-14B 0.800 0.614 9182 0.633 0.429 10046
Light-R1-14B-DS 0.833 0.641 9571 0.767 0.505 10194
LEAD-stage1 0.833 0.629 8790 0.767 0.523 9371
LEAD-stage3 0.867 0.650 8267 0.767 0.539 8668
5 Results
In this section, we present a comprehensive evaluation of the proposed GRPO-LEAD framework on two mathematical benchmarks: AIME24 and AIME25. Our analysis is structured as follows: we first examine training dynamics to illustrate how GRPO-LEAD accelerates convergence; next, we perform an ablation study to assess the incremental benefits of each component; and finally, we compare against state-of-the-art baselines for 14B-scale language models.
5.1 Training Dynamics
Figure 1 plots the evolution of Pass@1 on a validation split over training steps for three configurations of the 7B model: (i) baseline GRPO, (ii) GRPO with length reward, and (iii) GRPO with both length reward and advantage reweighting. We observe two clear trends. First, adding a length-dependent reward not only yields higher Pass@1 but also accelerates early-stage convergence, suggesting that penalizing overly verbose correct solutions provides a more informative learning signal. Second, incorporating advantage reweighting (to amplify updates on harder questions) further steepens the trajectory, indicating that reweighting advantage estimates according to problem difficulty helps the model refine reasoning on challenging prompts more efficiently.
Overall, these dynamics confirm that GRPO-LEAD components—particularly the length reward—bolster training stability and speed. By comparison, the baseline GRPO model learns more slowly and lags behind in Pass@1 across the entire training horizon.
5.2 Ablation Analysis
We next quantify the contribution of each GRPO-LEAD component through a step-by-step ablation on the 7B model. Table 1 summarizes results on AIME24 and AIME25.
Effect of Length Reward
We first incorporate the length-dependent accuracy reward into GRPO. Compared to Deepseek-7B, on AIME24, this maintains the original Cons@32 (0.767) while slightly improving Pass@1 from 0.431 to 0.438. Notably, the average solution length substantially decreases from 6990 to 5275 tokens, a reduction of approximately 24.5%. Similarly, on AIME25, Cons@32 improves from 0.467 to 0.533, accompanied by a significant length reduction of nearly 1900 tokens (approximately 26.8%). These results demonstrate, length reward, by penalizing correct but overly verbose solutions, can effectively reduce unnecessary text without compromising overall performance.
Effect of Advantage Reweighting
Adding difficulty-aware advantage reweighting further refines performance. On AIME24, although Cons@32 remains 0.767, Pass@1 increases to 0.458. On AIME25, both Cons@32 and Pass@1 improve (0.533 $\rightarrow$ 0.567 and 0.308 $\rightarrow$ 0.325, respectively). These results demonstrate that prioritizing challenging problems strengthens the model’s reasoning robustness, as the reweighting strategy mitigates over-reliance on simpler examples. This validates our hypothesis that calibrating training focus toward harder instances drives more reliable generalization.
Effect of Explicit Penalty for Incorrect Answers
Finally, we introduce a negative reward term to penalize incorrect solutions explicitly. As shown in Table 1, this addition yields the highest Pass@1 scores across the board (0.470 on AIME24 and 0.345 on AIME25). Cons@32 also climbs to 0.800 on AIME24. Note, however, that solution length increases modestly from about 5300 tokens to 6104 on AIME24, reflecting a trade-off: while the explicit penalty effectively sharpens the decision boundary and boosts accuracy, the model also tends to be more conservative and invests more tokens to ensure correctness. Nonetheless, the resulting average solution length is still lower than the baseline Deepseek-7B.
Overall, these ablation results confirm that all three enhancements—length-dependent accuracy, difficulty-aware advantage reweighting, and explicit penalties—collectively reduce verbosity, strengthen mathematical skills on harder questions, and elevate precision in final predictions.
5.3 Comparison with Baselines
We next evaluate GRPO-LEAD at the 14B scale and compare it against two strong baselines under a 14k-token generation budget: DeepSeek-14B and the state-of-the-art Light-R1-14B-DS. Table 2 presents results on AIME24 and AIME25, including both our intermediate model (LEAD-stage1) and our final model (LEAD-stage3).
AIME24 Performance
LEAD-stage1 achieves a Cons@32 of 0.833, matching Light-R1-14B-DS and exceeding DeepSeek-14B by 4.1% (0.833 vs. 0.800). Its Pass@1 (0.629) outperforms DeepSeek-14B (0.614) and closely approaches Light-R1-14B-DS (0.641). Crucially, LEAD-stage1 produces more concise responses (8790 tokens) than both baselines. Building on these gains, LEAD-stage3 pushes performance further, delivering the highest Cons@32 (0.867, 4% above Light-R1-14B-DS) and the best Pass@1 (0.650), while reducing average solution length to 8267 tokens.
AIME25 Performance
LEAD-stage1 yields a Cons@32 of 0.767, on par with Light-R1-14B-DS (0.767) and substantially ahead of DeepSeek-14B (0.633). Its Pass@1 (0.523) outperforms both DeepSeek-14B (0.429) and Light-R1-14B-DS (0.505). Once again, solutions from LEAD-stage1 are notably shorter (9371 tokens) than those of the baselines. In turn, LEAD-stage3 attains the highest Cons@32 (0.767) and Pass@1 (0.539), while further trimming average length to 8668 tokens.
Overall, both LEAD-stage1 and LEAD-stage3 deliver substantial improvements over DeepSeek-14B and Light-R1-14B-DS, simultaneously boosting correctness and conciseness under a constrained (14k-token) budget. Remarkably, training LEAD-stage1 for just 100 steps—requiring only about 24 hours on eight H20 GPUs—already matches Light-R1-14B-DS on Cons@32 and outperforms it on AIME25 Pass@1 while producing shorter solutions, underscoring the practical efficiency of GRPO-LEAD for large-scale math problem-solving.
6 Limitations
Although our techniques for encouraging concise solutions and difficulty-balanced learning may transfer to other domains, the gains reported here are specific to mathematical reasoning tasks. Further studies are needed to evaluate the effectiveness of GRPO-LEAD on broader question-answering or logical reasoning domains, where correctness signals and domain structures can differ substantially.
Additionally, we only have access to a limited amount of compute, which prevents us from conducting more comprehensive experiments. For instance, we currently cannot provide the validation curve for the 7B model in the ablation study that adds an explicit penalty. This is due to an error in the validation code after upgrading to the newest VERL version, and we currently don’t have the compute to reproduce it. We also couldn’t formally perform a hyperparameter search to showcase the rationale behind choosing the hyperparameters for our designed modifications.
7 Conclusion
We introduced GRPO-LEAD, a reinforcement learning framework designed for mathematical reasoning tasks. By extending Group Relative Policy Optimization with three major components—(1) a length-dependent accuracy reward to discourage overly verbose solutions, (2) an explicit negative penalty that clarifies the boundary between correct and incorrect answers, and (3) a difficulty-aware advantage reweighting scheme to prioritize tougher problems—GRPO-LEAD addresses key challenges in structured problem-solving.
Empirical evaluations on two AIME benchmarks show that GRPO-LEAD not only speeds up convergence but also strengthens the model’s reasoning capability while keeping solution paths concise. Our 14B-scale experiments further confirm that GRPO-LEAD achieves state-of-the-art performance by balancing output brevity with high problem-solving accuracy. Although open questions remain—particularly in managing partial correctness and extending these techniques to broader domains—our findings suggest that reward shaping and difficulty modeling are pivotal in developing more robust and aligned language models for complex mathematical reasoning.